# Revealing the Origin of Neutrino Masses through Displaced Shower Searches in the CMS Muon System

### v1

@@ III Displaced Showers in the CMS Muon System
[…]
@@
We determine the sensitivity of the CMS search Tumasyan et al. (2021), which uses the muon detector as a sampling calorimeter to identify showers produced by the decays of an LLP. An update to this search is available Hayrapetyan et al. (2024b). However, detailed reinterpretation information is, as of yet, missing. We do not expect this to have a large impact as the triggers are largely the same. The decays produce photons, electrons, tau or quarks inside the muon detector creating hadronic and electromagnetic showers and giving rise to a large number of hits in a small region of the detector as the showers are collimated. These hits are identified by cathode strip chambers (CSCs), which detect charged particles via ionisation of gas molecules and they are clustered by identifying high-density regions. Such clusters are called CSC clusters. Using this strategy, the CMS search detects electromagnetic and hadronic energy from electrons, taus, photons and jets but it does not include muons as they rarely produce particle showers. As the search strategy relies on identification and reconstruction of original information using CSC clusters, to aid reinterpretation of the search, CMS provides an updated card and modules for the Delphes detector simulation Wang (2022); CMS Collaboration (2021); Tumasyan et al. (2021). We implement these to determine the reconstruction efficiency and kinematic information of the CSC cluster events, and apply the selection criteria of the CMS search:
@@
The missing transverse energy is $E_{T}^{\text{miss}}>200$ GeV as trigger requirement. The missing transverse energy is defined as the negative vector sum of visible $p_{T}$ energy deposited in the tracker and calorimeter. If the signal contains only LLPs which decay beyond the calorimeter, there is no missing energy to trigger on. We thus simulate the signal up to two initial state jets, which are prompt and deposit energy in the calorimeter.
@@
No electron (muon) with transverse momentum $p_{T}>35~{}(25)$ GeV and pseudorapidity $|\eta|<2.5~{}(2.4)$, to remove background from $W$ and top production.
@@
At least one CSC cluster with $|\Delta\phi(\mathbf{x}_{\text{CSC}},\mathbf{p}_{T}^{\text{miss}})|<0.75$ to ensure that the cluster originated from the LLP decay. Here, $\Delta\phi$ is defined as the azimuthal angle between the missing transverse momentum and the cluster location from the IP. This ensures that the LLP is close to the direction of missing energy, which points in the opposite direction of the vector sum of the visible $p_{T}$.
@@
Events with clusters too close to a jet (muon) are removed, for $\Delta R=\sqrt{(\Delta\eta)^{2}+(\Delta\phi)^{2}}<0.4$. This ensures that the CSC cluster is not matched to jets (muons) with $p_{T}>10~{}(20)$ GeV and the it is not created by LLPs inside the jet, e.g. $K_{L}$, or muon bremsstrahlung.
@@
The average time of detector hits in the CSC cluster, relative to the time of collision is $-5~{}\text{ns}<\langle\Delta t_{\text{CSC}}\rangle<12.5$ ns, to reject clusters produced by pileup.
@@
Figure 1: Missing transverse energy $E_{T}^{\text{miss}}$ for $pp\to Z^{\prime}\to NN$ ($m_{Z^{\prime}}=500$ GeV, blue), $pp\to Z\to NN$ (green), $pp\to h\to NN$ (red) and $pp\to\Phi\to NN$ ($m_{\Phi}=150$ GeV, orange), with $m_{N}=30$ GeV and $|V_{\ell N}|^{2}=10^{-12}$. The vertical dashed lines indicate the thresholds $E_{T}^{\text{miss}}>200$ GeV and $>50$ GeV in the CMS search and the soft trigger strategy, respectively.
@@
Due to $E_{T}^{\text{miss}}>200$ GeV, only a small number of signal events passes the selection criteria. This is illustrated in Fig. 1, showing the corresponding distribution for our production modes. The rate drops rapidly for larger $E_{T}^{\text{miss}}$, especially for SM $Z$ production, due to the fact that the missing transverse energy either originates from initial state radiation, which is enhanced for Higgs gluon-gluon fusion, or from heavier resonant particle masses, as in the $Z^{\prime}$ and $\Phi$ modes. Nevertheless, only $\approx 1\%$ of events satisfy the trigger requirement. In the SM $Z$ mode, only $\lesssim 0.1\%$ pass the trigger and we thus expect no appreciable sensitivity in this case.
@@
To calculate the dominant gluon-gluon fusion Higgs production, we use the effective gluon-gluon-Higgs coupling at leading order Alwall et al. (2011). We have checked that this is accurate for our purposes, giving a conservative determination; while the transverse momentum of the Higgs can significantly change when including higher order corrections Jones et al. (2018); Becker et al. (2024), by comparing the leading order effective treatment with the next-to-leading order fully resolved top quark mass-dependent vertex model Maltoni et al. (2014), the effective model leads to a softer Higgs for $p_{T}<200$ GeV, while the distributions are comparable for $400200<p_{T}<$ 400 GeV. For $p_{T}>400$ GeV, the effective model overestimates by more than a factor of two. Nevertheless, as its cross section is roughly an order of magnitude smaller than for $400200<p_{T}<$ 400 GeV, we consider our calculation accurate for $p_{T}>200$ GeV.
@@
In Ref. Cottin et al. (2023), it is argued that the CMS search could be improved as a search for sterile RHNs by assuming a dedicated Level-1 and a high-level displaced signature trigger. The requirement on $E_{T}^{\text{miss}}$ can then be relaxed to $E_{T}^{\text{miss}}>50$ GeV, resulting in an improved signal rate by about an order of magnitude, cf. Fig. 1, while the background can still be controlled by requiring larger CSC clusters. We will refer to this option as the ‘soft trigger strategy’ in the following. The softened $E_{T}^{\text{miss}}$ threshold means that we are more susceptible to Higgs production corrections as we are probing softer Higgs $p_{T}$. Our results are conservative as we underestimate the Higgs $p_{T}$ distribution by about a factor of two.
@@
We simulate the displaced shower signature at the 13 TeV LHC with 137 fb-1 of integrated luminosity and the 14 TeV HL-LHC with 3000 fb-1. The background mainly comes from punch-through jets and muon bremsstrahlung, controlled by the above selection criteria and determined from CMS data in Tumasyan et al. (2021): The number of expected background events is $b=2.0\pm 1.0$ and the number of observed events is $N=3$. excluding an expected number of signal events of $s>6.13$ at 95% confidence level (CL). At the HL-LHC, after taking into account a 20% (60%) signal loss due to pileup, and requiring more hits to control the background, $s>56~{}(3)$ is required at 95 % CL with the CMS (soft) trigger strategy Cottin et al. (2023). For the HL-LHC projection, we assume $b=40$ background events at the HL-LHC for the CMS trigger and no background for the soft trigger.
@@ IV Probing the Neutrino Mass Generation Mechanism
Using the above simulation strategy, we reinterpret the CMS search Tumasyan et al. (2021) in the $B-L$ model, specifically to constrain or probe the RHN mass $m_{N}$ and the active-sterile mixing $|V_{\ell N}|$ for successful neutrino mass generation. In our simplified model with a single RHN $N$, a light neutrino acquires a Majorana mass $m_{\nu}=|V_{\ell N}|^{2}m_{N}$. The KATRIN experiment limits the effective $\beta$ decay neutrino mass as $m_{\beta}<0.45$ eV at 90 % CL Aker et al. (2024), whereas solar neutrino oscillations point towards a neutrino mass scale of $\sqrt{\Delta m^{2}_{\text{sol}}}=9\times 10^{-3}$ eV Esteban et al. (2020). We take the range $0.459\times 10^{-3}~{}\text{eV}<|V_{\ell N}|^{2}m_{N}<$ 0.45 eV as a target to probe the canonical seesaw floor of neutrino mass generation. We only consider active-sterile mixing to $\ell=e,\tau$ flavour since the CMS muon detector cannot respond to displaced muons.
@@ IV.1 Gauge Portal
We start with the $Z^{\prime}$ gauge portal $pp\to Z^{\prime}\to NN$ where we use the benchmark values $m_{Z^{\prime}}=500$ GeV and $g_{B-L}=0.005$, satisfying current constraints. No sensitivity to the neutrino mass generation parameters is obtained in this case using the current CMS data with either the original or the soft trigger strategy, since the strong requirement on $E_{T}^{\text{miss}}$ cuts away most of the signal. At the HL-LHC with 3000 fb-1, our reinterpretation also only predicts a sensitivity for the soft trigger strategy. If the $B-L$ gauge coupling is larger than $g_{B-L}\gtrsim 0.0035$, RHNs can be probed in this case, within the mass range $10~{}\text{GeV}\lesssim m_{N}\lesssim 80$ GeV.
Prospects for the SM gauge portal $pp\to Z\to NN$ are similar. For a $ZZ^{\prime}$ mixing $\beta=10^{-3}$ at the current upper limit, no sensitivity at 95 % CL is achieved using current data. It can only be probed at the HL-LHC using the soft trigger strategy, for $m_{N}\lesssim 40$ GeV. While the total cross section for the SM $Z$ is sizeable as discussed in Sec. II, Fig. 1 demonstrates that a large number of events from $Z$ production are removed due to small $E_{T}^{\text{miss}}$.
@@ IV.2 Higgs Portal
Figure 2: Sensitivity of $pp\to h\to NN$ (left) and $pp\to\Phi\to NN$ (right, $m_{\Phi}=150$ GeV) at 95% CL on the RHN mass $m_{N}$ and the active-sterile mixing $|V_{\ell N}|^{2}$ ($\ell=e,\tau$) using displaced shower searches in the CMS muon system. The scenarios are for the reinterpretation of the CMS data (red region), as well as the 14 TeV HL-LHC with 3000 fb-1 using the CMS (blue contours) and soft (red contours) trigger strategies. The Higgs mixing is fixed at $\sin\alpha=0.24$ (red region and solid contours) and $0.14$ (dashed). The $B-L$ Higgs VEV is fixed at $\langle\Phi\rangle=3.75$ TeV. The region labelled ’Current Limits’ is excluded by existing sterile neutrino searches ($\ell=e$) while the contours labelled SHiP, CMS and FCC-ee give the projected sensitivity of planned experiments to sterile neutrinos, compiled in Bolton et al. (2020, 2022). The band labelled ‘Seesaw’ indicates where $0.459\times 10^{-3}~{}\text{eV}<|V_{\ell N}|^{2}m_{N}<$ 0.45 eV.
@@
The prospects for the Higgs portals are much more promising. Applying the above analysis to the SM Higgs case, $pp\to h\to NN$, reinterpreting the CMS search, we find that existing CMS data already constrains the parameter space for light neutrino mass generation in the $B-L$ gauge model. This is shown in Fig. 2 (left), where the red region labelled ‘CMS’ is excluded at 95% CL, for a Higgs mixing $\sin\alpha=0.24$ and $B-L$ VEV $\langle\Phi\rangle=3.75$ TeV. As anticipated in Sec. II.2, the sensitivity is maximal for $m_{N}\approx m_{h}/\sqrt{10}\approx 40$ GeV. For a displaced search at distances $\mathcal{O}(1~{}\text{m})$ in the CMS muon system, this corresponds to the well-motivated canonical seesaw prediction $|V_{\ell N}|^{2}\approx 10^{-12}$ that is being probed.
@@
Fig. 2 (left) also shows the projected sensitivity at the HL-LHC using the CMS (solid blue contour) and soft trigger (solid red contour) strategies. This extends the coverage of the seesaw regime over a broader range, up to $10~{}\text{GeV}\lesssim m_{N}\lesssim 60$ GeV. The HL-LHC will improve the sensitivity to the Higgs mixing, to values $\sin\alpha\gtrsim 0.14$ Cepeda et al. (2019); ATLAS and CMS (2019). Using this value, i.e., assuming the HL-LHC will see no sign of the Higgs mixing with an exotic state, the sensitivity decreases to the dashed contours in Fig. 2 (left). In the three scenarios thus considered, the smallest $\sin\alpha$ for which part of the seesaw region can be probed at 95% CL is $\sin\alpha>0.17$ (CMS reinterpretation), $0.1$ (HL-LHC with CMS trigger) and $0.017$ (HL-LHC with soft trigger).
@@
In a broader context, beyond the scope of the $B-L$ model, this result can also be interpreted in terms of the operator $\mathcal{O}_{NH}=(\overline{N^{c}}N)(H^{\dagger}H)/\Lambda_{\text{NP}}$ in the RHN-extended SM effective field theory Barducci et al. (2021), assuming that there are no other exotic light degrees of freedom other than the RHN. After EW symmetry breaking it leads to the vertex $(v/\Lambda_{\text{NP}})hNN$, which will allow the SM Higgs to decay to RHNs. Matching the vertex factor with that in the $B-L$ model, $v/\Lambda_{\text{NP}}=y_{N}\sin\alpha/2=m_{N}\sin\alpha/(2\langle\Phi\rangle)$, we can recast the above sensitivities to the New Physics operator scale $\Lambda_{\text{NP}}$: $\Lambda_{\text{NP}}\approx 400$ TeV (CMS reinterpretation), $\Lambda_{\text{NP}}\approx 600$ TeV (HL-LHC with CMS trigger) and $\Lambda_{\text{NP}}\approx 4000$ TeV (HL-LHC with soft trigger), applicable for a RHN with $m_{N}=40$ GeV and $|V_{\ell N}|^{2}=10^{-12}$. These are already comparable to prospects at future lepton colliders Barducci et al. (2021).
@@
As discussed in Sec. II.2, the exotic Higgs production channel $pp\to\Phi\to NN$ can have a large cross section and is nearly unsuppressed by the Higgs mixing $\sin\alpha$ for $m_{h}<m_{\Phi}\lesssim 150$ GeV. Consequently, the current constraint and the future sensitivity are even better than for SM Higgs production, as shown in Fig. 2 (right). The same parameters and scenarios are considered as above, with an exotic Higgs mass of $m_{\Phi}=150$ GeV. As can be seen, a reduction of the Higgs mixing from $\sin\alpha=0.24$ to $0.14$ has little impact due to the decreased sensitivity on the mixing. Hence, the weakest Higgs mixing that still results in a 95% CL signal can be as small as $\sin\alpha=0.07$ (CMS reinterpretation), $0.02$ (HL-LHC with CMS trigger) and $0.002$ (HL-LHC with soft trigger).

### v2

@@ III Displaced Showers in the CMS Muon System
The overall rates above motivate a detailed simulation to match the CMS endcap search. The model is implemented in Universal FeynRules Output (UFO) [88], following [64, 53]. The event generator MadGraph5aMC@NLO v3.4.1 [89] is then used for a parton level simulation of the events with the shower jets and matrix element jets matched [90, 91]. Events are fed to PYTHIA v8.235 [92] for parton showering, hadronization and heavy hadron decays. Clustering of events is performed by FastJet v3.2.1 [93] and detector effects by Delphes v3.5.1 [94].
@@
The CMS search [75], which uses the muon detector as a sampling calorimeter, identifies LLP showers. The updated search [95] might improve the detector efficiency but there is no detailed reinterpretation information available yet. There is also a similar search for vector-like leptons [96]. We do not expect this to have a large impact as the triggers are largely the same. LLP decays produce photons, electrons, tau or quarks inside the muon detector creating collimated hadronic and electromagnetic showers and giving rise to a large number of hits in a small detector region. These hits are identified by cathode strip chambers (CSCs), which detect charged particles and they are clustered (called CSC clusters) by identifying high-density regions. As the search strategy relies on identification and reconstruction of original information using CSC clusters, to aid reinterpretation, CMS provides an updated card and modules for Delphes [97, 98, 75]. We implement these to determine the reconstruction efficiency and kinematic information of the CSC cluster events, and apply the selection criteria of the CMS search:
@@
The missing transverse energy, defined as the negative vector sum of visible $p_{T}$ energy deposited in the tracker and calorimeter, is $E_{T}^{\text{miss}}>200$ GeV, as trigger requirement. If the signal contained only LLPs decaying beyond the calorimeter, there is nothing to trigger on. We thus simulate the signal up to two initial state jets, which are prompt and deposit energy in the calorimeter.
@@
No electron (muon) with transverse momentum $p_{T}>35~{}(25)$ GeV and pseudorapidity $|\eta|<2.5~{}(2.4)$, to remove $W$ and top background.
@@
At least one CSC cluster with $|\Delta\phi|<0.75$ to ensure that it originated from the LLP decay. Here, $\Delta\phi(\mathbf{x}_{\text{CSC}},\mathbf{p}_{T}^{\text{miss}})$ is defined as the azimuthal angle between the missing transverse momentum and the cluster location from the IP. The LLP is then close to the missing momentum, which points opposite to the vector sum of the visible $p_{T}$.
@@
Events with clusters too close to a jet (muon) are removed, for $\Delta R=\sqrt{(\Delta\eta)^{2}+(\Delta\phi)^{2}}<0.4$. The CSC cluster is then not matched to jets (muons) with $p_{T}>10~{}(20)$ GeV and it is not created by LLPs inside the jet, e.g. $K_{L}$, or muon bremsstrahlung.
@@
The average time of detector hits in the CSC cluster, relative to the collision is $-5~{}\text{ns}<\langle\Delta t_{\text{CSC}}\rangle<12.5$ ns, to reject pileup clusters.
@@
Figure 1: Missing transverse energy $E_{T}^{\text{miss}}$ for $pp\to Z^{\prime}\to NN$ ($m_{Z^{\prime}}=500$ GeV, blue), $Z$ (green), $h$ (red) and $\Phi$ ($m_{\Phi}=150$ GeV, orange), with $m_{N}=30$ GeV and $|V_{\ell N}|^{2}=10^{-12}$. The vertical lines indicate the thresholds $E_{T}^{\text{miss}}>200$ GeV and $>50$ GeV in the CMS search and the soft trigger strategy, respectively.
@@
Due to $E_{T}^{\text{miss}}>200$ GeV, only a small number of events passes the selection criteria, see Fig. 1. The rate drops rapidly for larger $E_{T}^{\text{miss}}$, especially for SM $Z$ production, due to the fact that missing transverse energy either originates from initial state radiation, which is enhanced for Higgs gluon-gluon fusion, or from heavier resonant particle masses, as in the $Z^{\prime}$ and $\Phi$ modes. Nevertheless, only $\approx 1\%$ of events satisfy the trigger requirement. In the SM $Z$ mode, only $\lesssim 0.1\%$ pass the trigger and we thus expect no appreciable sensitivity in this case.
@@
To calculate the dominant gluon-gluon fusion Higgs production, we use the effective gluon-gluon-Higgs coupling at leading order [99]. We have checked that this is accurate, giving a conservative determination; while the transverse momentum of the Higgs can change when including higher order corrections [100, 101], by comparing the leading order effective treatment with the next-to-leading order fully resolved top quark mass-dependent vertex model [102], the effective model leads to a softer Higgs for $p_{T}<200$ GeV, while the distributions are comparable for $400200~{}\text{GeV}<p_{T}<$ 400 GeV. For $p_{T}>400$ GeV, the effective model overestimates by more than a factor of two. Nevertheless, as the cross section is much smaller than for $400200~{}\text{GeV}<p_{T}<$ 400 GeV, we consider our calculation accurate for $p_{T}>200$ GeV.
@@
Given that new, dedicated Level-1 and high level triggers targeting our signature have been collecting data during the LHC Run-3 [103], Ref. [79] argues that the requirement on $E_{T}^{\text{miss}}$ can be relaxed to $E_{T}^{\text{miss}}>50$ GeV, resulting in an improved signal rate, cf. Fig. 1, while the background can still be controlled by requiring larger CSC clusters. We refer to this option as the ‘soft trigger strategy’. The softened $E_{T}^{\text{miss}}$ threshold means that we are more susceptible to Higgs production corrections as we are probing softer Higgs $p_{T}$. Our results remain conservative as we underestimate the Higgs $p_{T}$ distribution by about a factor of two.
@@
We simulate the displaced shower signature at the 13 TeV LHC with 137 fb-1 integrated luminosity and the 14 TeV HL-LHC with 3000 fb-1. The background mainly comes from punch-through jets and muon bremsstrahlung, controlled by the above selection criteria and determined from CMS data: The number of expected background events is $b=2.0\pm 1.0$ and the number of observed events is $N=3$, excluding signal events $s>6.1$ at 95% CL [75]. At the HL-LHC, after taking into account a 20% (60%) signal loss due to pileup, and requiring more hits to push the background to be negligible, $s>56~{}(3.0)$ is required at 95 % CL with the CMS (soft) trigger strategy, where we scale the number of background events according to luminosity, $b\approx 40~{}(0.0)$ at the HL-LHC with the CMS (soft) trigger [79].
@@ IV Probing the Neutrino Mass Generation Mechanism
Using the above strategy, we reinterpret the CMS search [75] in the $B-L$ model and the $\nu_{R}$, specifically to probe the RHN mass $m_{N}$ and the active-sterile mixing $|V_{\ell N}|$ for successful neutrino mass generation. In our scenarios, with a single RHN $N$, a light neutrino acquires the Majorana mass $m_{\nu}=|V_{\ell N}|^{2}m_{N}$. As indicated in the introduction, we take the range $0.459\times 10^{-3}~{}\text{eV}<|V_{\ell N}|^{2}m_{N}<$ 0.45 eV as a target to probe the seesaw floor. We only consider active-sterile mixing to $\ell=e,\tau$ flavour as mentioned.
@@ IV.1 Gauge Portal
We start with the $Z^{\prime}$ gauge portal $pp\to Z^{\prime}\to NN$ where we use the benchmark values $m_{Z^{\prime}}=500$ GeV and $g_{B-L}=5\times 10^{-3}$. No appreciable sensitivity is obtained in this case using the CMS data with either the original or the soft trigger strategy, since the requirement on $E_{T}^{\text{miss}}$ removes most of the signal. At the HL-LHC with 3000 fb-1, our reinterpretation also only predicts a sensitivity for the soft trigger strategy. If the $B-L$ gauge coupling is larger than $g_{B-L}\gtrsim 0.0035$, RHNs can be probed in this case, within the mass range $10~{}\text{GeV}\lesssim m_{N}\lesssim 80$ GeV.
Prospects for the SM gauge portal $pp\to Z\to NN$ are similar. For a $ZZ^{\prime}$ mixing $\beta=10^{-3}$, no sensitivity at 95% CL is achieved using current data. It can only be probed at the HL-LHC using the soft trigger strategy, for $m_{N}\lesssim 40$ GeV. While the total cross section for $Z$ is sizeable, Fig. 1 demonstrates that a large number of events are removed due to small $E_{T}^{\text{miss}}$.
@@ IV.2 Higgs Portal
Figure 2: Sensitivity of $pp\to h\to NN$ at 95% CL on the RHN mass $m_{N}$ and the active-sterile mixing $|V_{\ell N}|^{2}$ ($\ell=e,\tau$). The scenarios are for reinterpreting CMS data (red region), as well as at the 14 TeV HL-LHC with 3000 fb-1 using the CMS (blue contour) and soft (green contour) trigger strategies. The Higgs mixing is $\sin\alpha=0.24$ (red region and solid contours) and $0.14$ (dashed), and the $B-L$ Higgs VEV is $\langle\Phi\rangle=3.75$ TeV. The region labelled ‘Current Limits’ is excluded by sterile neutrino searches ($\ell=e$) while the contours labelled SHiP, CMS and FCC-ee give the projected sensitivity of planned experiments [12, 104]. The band labelled ‘Seesaw’ indicates $0.459\times 10^{-3}~{}\text{eV}<|V_{\ell N}|^{2}m_{N}<$ 0.45 eV.
@@
Prospects for the Higgs portals are much more promising. Reinterpreting the CMS search for $pp\to h\to NN$, we find that existing CMS data constrains parameter space of light neutrino mass generation in the $B-L$ model. This is shown in Fig. 2, where the red region labelled ‘CMS’ is excluded at 95% CL, for a Higgs mixing $\sin\alpha=0.24$ and $B-L$ VEV $\langle\Phi\rangle=3.75$ TeV. As anticipated in Sec. II.2, the sensitivity is maximal for $m_{N}\approx m_{h}/\sqrt{10}\approx 40$ GeV. For a displaced search at distances $\mathcal{O}(1~{}\text{m})$ in the CMS muon system, this corresponds to the well-motivated seesaw prediction $|V_{\ell N}|^{2}\approx 10^{-12}$.
@@
Fig. 2  also shows the projected sensitivity at the HL-LHC using the CMS (solid blue contour) and soft trigger (solid red contour) strategies. This extends the coverage over $10~{}\text{GeV}\lesssim m_{N}\lesssim 60$ GeV. The HL-LHC will improve the sensitivity to the Higgs mixing, to values $\sin\alpha\gtrsim 0.14$ [105, 106]. Using this value, i.e., assuming the HL-LHC will see no sign of the Higgs mixing with an exotic state, the sensitivity decreases to the dashed contours in Fig. 2. In the three scenarios considered, the smallest $\sin\alpha$ for which part of the seesaw region can be probed at 95% CL is $\sin\alpha>0.17$ (CMS reinterpretation), $0.1$ (HL-LHC with CMS trigger) and $0.017$ (HL-LHC with soft trigger).
Figure 3: As Fig. 2, but showing the sensitivity of the CMS search to the scale $\Lambda_{\text{NP}}/|\mathcal{C}_{NH}|$ of the $\nu_{R}$ operator $\mathcal{O}_{NH}$.
@@
Our result can also be interpreted in terms of the $\nu_{R}$ operator $\mathcal{O}_{NH}=(\bar{N}^{c}N)(H^{\dagger}\cdot H)$. It is the RHN-equivalent of the Weinberg operator, appearing at the same mass dimension-5 and it leads to the vertex $(v\mathcal{C}_{NH}/\Lambda_{\text{NP}})hNN$, enabling the decay $h\to NN$. Here, $\Lambda_{\text{NP}}$ is the New Physics scale in which the $\nu_{R}$ is expanded and $\mathcal{C}_{NH}$ is the dimensionless Wilson coefficient of $\mathcal{O}_{NH}$, see Appendix B for a brief summary of the $\nu_{R}$. Matching the vertex factor with that in the $B-L$ model, $v\mathcal{C}_{NH}/\Lambda_{\text{NP}}=y_{N}\sin\alpha/2=m_{N}\sin\alpha/(2\langle\Phi\rangle)$, we can recast the above sensitivities to the $\mathcal{O}_{NH}$ operator scale. As can be seen in Fig. 3, operator scales as high as $\Lambda_{\text{NP}}/|\mathcal{C}_{NH}|\approx 300$ TeV are already being probed using the CMS data. At the HL-LHC using the CMS and soft trigger strategies, scales up to $\Lambda_{\text{NP}}/|\mathcal{C}_{NH}|\approx 500$ TeV and $2000$ TeV can be tested, respectively. These are already comparable to prospects at future lepton colliders [107].
@@
As discussed in Sec. II.2, the exotic Higgs channel $pp\to\Phi\to NN$ can have a large cross section and is nearly unsuppressed by the Higgs mixing $\sin\alpha$ for $m_{h}<m_{\Phi}\lesssim 150$ GeV. The current constraint and future sensitivities are qualitatively similar but better than for SM Higgs production. With the same parameters and scenarios considered above and an exotic Higgs mass of $m_{\Phi}=150$ GeV, the smallest Higgs mixing that still results in a 95% CL signal is $\sin\alpha\approx 0.07$ (CMS reinterpretation), $0.02$ (HL-LHC with CMS trigger) and $0.002$ (HL-LHC with soft trigger).